OpenStudy (lgbasallote):

another simple one but it was dead end for me..pls show me steps :DD \(\LARGE \int \frac{x^3}{\sqrt{x^2 + 4}} dx\) @apoorvk

OpenStudy (aravindg):

:P

OpenStudy (aravindg):

shall i try?

OpenStudy (lgbasallote):

trig substitution... @AravindG sure..

OpenStudy (aravindg):

try x=2tan theta

OpenStudy (lgbasallote):

i did...it brought me much later on to \(\LARGE \sec \theta \tan \theta (\sec^2 \theta -1) dx\)

OpenStudy (aravindg):